### v1

# Efficient Testing Implies Structured Symmetry

Abstract
Given a small random sample of $n$-bit strings labeled by an unknown Boolean function, which properties of this function can be tested computationally efficiently? We show an equivalence between properties that are efficiently testable from few samples and properties with structured symmetry, which depend only on the function’s average values on parts of a low-complexity partition of the domain. Without the efficiency constraint, a similar characterization in terms of unstructured symmetry was obtained by Blais and Yoshida (2019). Our main technical tool is supersimulation, which builds on methods from the algorithmic fairness literature to approximate arbitrarily complex functions by small-circuit simulators that fool significantly larger distinguishers.
We extend the characterization along other axes as well. We show that allowing parts to overlap exponentially reduces their required number, broadening the scope of the construction from properties testable with $O(\log n)$ samples to properties testable with $O(n)$ samples. For larger sample sizes, we show that any efficient tester is essentially checking for indistinguishability from a bounded collection of small circuits, in the spirit of a characterization of testable graph properties. Finally, we show that our results for Boolean function testing generalize to high-entropy distribution testing on arbitrary domains.
This work was supported in part by Simons Foundation Grant 733782 and Cooperative Agreement CB20ADR0160001 with the United States Census Bureau.
1 Introduction
The goal in distribution testing [GGR96] is to determine whether an unknown probability distribution $\mathcal{D}$ has a particular property of interest, such as uniformity, independence, or equality to a reference distribution. Rather than being given a complete description of $\mathcal{D}$, the algorithm has access to an oracle that generates independent samples from $\mathcal{D}$ on request. Ideally, we would like to make as few requests as possible, run a fairly simple computation on the resulting samples, and determine whether $\mathcal{D}$ has the property or is far from having the property.
In this work, we focus on distributions over pairs $(x,y)\in\{0,1\}^{n}\times\{0,1\}$, where $x$ is a uniform $n$-bit string, and $y=f(x)$ for some unknown Boolean function $f:\{0,1\}^{n}\to\{0,1\}$. Our main result is a characterization of which properties of $f$ can be efficiently tested from few samples in this framework. By “efficient,” we mean that the testing algorithm can be represented as a small Boolean circuit that takes as input a set of labeled samples and outputs Accept or Reject.
1.1 Background
Without the requirement of computational efficiency, an elegant result of Blais and Yoshida [BY19] fully characterizes the properties of Boolean functions that are testable from a constant number of samples (i.e. not scaling with $n$). They showed that a property is constant-sample testable if and only if it has constant-part symmetry, where a property of $f$ is said to have $k$-part symmetry if it depends only on the average values of $f$ within some fixed partition of the domain into parts $S_{1},\ldots,S_{k}\subseteq\{0,1\}^{n}$. Equivalently, permutations of the domain within each part do not affect whether or not $f$ has the property.
More formally, let $\mathcal{P}$ be a property, which we represent as a set of Boolean functions. Let $\mathcal{P}_{\varepsilon}$ denote the set of functions that are $\varepsilon$-close to some function in $\mathcal{P}$, where distance is measured by the fraction of inputs on which two functions disagree. We say that a tester has proximity parameter $\varepsilon$ if it accepts all $f\in\mathcal{P}$ with probability at least $2/3$ and rejects all $f\notin\mathcal{P}_{\varepsilon}$ with probability at least $2/3$ (see Section 2 for more detail on the setup). In this language, Blais and Yoshida showed:
Theorem 1.1 ([BY19]).
If a property $\mathcal{P}$ of Boolean functions is testable with proximity $\varepsilon$ using $m$ samples, then $\mathcal{P}\subseteq\mathcal{Q}\subseteq\mathcal{P}_{\varepsilon}$ for some $2^{2^{O(m)}}$-part symmetric property $\mathcal{Q}$. [^2] Note that $m$ may scale with $n$ and $\varepsilon$ arbitrarily, but the conclusion will be vacuous if $2^{2^{O(m)}}$ exceeds $2^{n}$. Conversely, any $k$-part symmetric property $\mathcal{Q}$ is testable with proximity $\varepsilon$ using $(k/\varepsilon)^{O(1)}$ samples.
Theorem 1.1 gives a transformation from $\mathcal{P}$ to $\mathcal{Q}$ that loses computational efficiency (ours will not). Indeed, $\mathcal{Q}$ may be computationally intractable even when $\mathcal{P}$ has a computationally efficient tester. To illustrate this, consider any $k$-part symmetric property $\mathcal{Q}$ with invariant sets $S_{1},\ldots,S_{k}$. Clearly, $\mathcal{Q}$ can be tested from a handful of $(x_{i},y_{i})$ samples—just classify each $x_{i}$ according to the part $S_{j}$ that contains it, and use the corresponding labels $y_{i}$ to empirically estimate the average value of $f$ over $S_{j}$. Given these $k$ averages, testing for $\mathcal{Q}$ can always be done efficiently, simply because any function on few inputs can be computed with a small circuit by brute force. However, the overall procedure may still be computationally inefficient—indeed, because the sets $S_{j}$ are unstructured, performing even a single classification could require up to $\exp(n)$ time!
1.2 Results Overview
The previous example raises a natural question: does there exist an analogue of Theorem 1.1 that characterizes the properties of Boolean functions that are efficiently testable from a constant number of samples? Our first result shows that this is indeed the case. We prove that if the tester for $\mathcal{P}$ is a small circuit, then $\mathcal{P}$ is close to a property $\mathcal{Q}$ with structured symmetry. By this, we mean that $\mathcal{Q}$ is not only $k$-part symmetric with respect to some partition $S_{1},\ldots,S_{k}$, but also has a classifier circuit of size at most $s$ that computes the index $j\in[k]$ of the part $S_{j}$ containing a given input $x\in\{0,1\}^{n}$. In this case, we say that $\mathcal{Q}$ has computational partition complexity $s$, or simply partition complexity $s$.
Theorem 1.2.
If $\mathcal{P}$ is testable with proximity $\varepsilon$ using $m$ samples and a circuit of size $s$, then $\mathcal{P}\subseteq\mathcal{Q}\subseteq\mathcal{P}_{\varepsilon}$ for some $2^{2^{O(m)}}$-part symmetric property $\mathcal{Q}$ with partition complexity $2^{O(m)}s$.
Theorem 1.3.
Any $k$-part symmetric property $\mathcal{Q}$ with partition complexity at most $s$ is testable with proximity $\varepsilon$ using $(k/\varepsilon)^{O(1)}$ samples and a circuit of size $(k/\varepsilon)^{O(1)}s+(k/\varepsilon)^{O(k)}$.
Consider the case that $m$, $k$, and $\varepsilon$ are constants, so that the circuit size $s$ is the only parameter varying with $n$. In this case, Theorem 1.1 states that constant-sample testability is equivalent to constant-part symmetry, while Theorems 1.2 and 1.3 state that constant-sample testability with a circuit of size $O(s)$ is equivalent to constant-part symmetry with partition complexity $O(s)$.
To prove Theorems 1.2 and 1.3, our main technical tool is supersimulation [DT25], which builds on a series of constructive generalizations of the complexity-theoretic regularity lemma [TTV09] pioneered by [HKRR18] in the algorithmic fairness literature (see also [DKR+21, GKR+22, CDV24], which are more closely related). The basic versions of these results take labeled samples of the form $(x_{i},f(x_{i}))$ and a function class $\mathcal{F}$ as input, and build a simulator $\tilde{f}$ making few oracle calls to members of $\mathcal{F}$ and fooling distinguishers in $\mathcal{F}$. Supersimulators, however, also fool distinguishers that are significantly more powerful than themselves.
Our proof first simulates the oracle that generates the samples, and then simulates the tester itself. Interestingly, between the two steps, the roles of the simulators and distinguishers switch. We discuss our proof strategy further in the technical overview in Section 1.3, and give a full proof in Section 3.
Extensions
In Section 4, we establish various other results regarding property testing of Boolean functions from samples, beyond our main results (Theorems 1.2 and 1.3).
First, in Section 4.1, we address another drawback of the result of [BY19]: the doubly exponential dependence on the tester’s sample complexity $m$ in the number of parts of the partition $S_{1},\ldots,S_{k}$. We show that allowing parts to overlap can exponentially improve this dependence, reducing the number of structured sets from $2^{2^{O(m)}}$ to just $2^{O(m)}$. We actually show an even stronger result: If $\mathcal{P}$ is $O(m)$-sample testable, then there exists a tester for $\mathcal{P}$ which simply counts the number of functions $f\in\mathcal{F}_{\mathrm{good}}$ and $f\in\mathcal{F}_{\mathrm{bad}}$ that are consistent with its $O(m)$ observed samples—the tester outputs Accept if these samples are consistent with more good functions than bad, and Reject otherwise. Here, $\mathcal{F}_{\mathrm{good}}$ and $\mathcal{F}_{\mathrm{bad}}$ are some prespecified families of at most $2^{O(m)}$ Boolean functions, which depend on the property $\mathcal{P}$. For this result, we do not consider their computational complexity. From another point of view, a $k$-part partition of a domain of size $2^{n}$ is nontrivial only if $k<2^{n}$. Thus, constructing an $2^{2^{O(m)}}$-part partition is only interesting for properties testable with $m=O(\log n)$ samples, but our theorem remains meaningful for a broader class of properties, namely those testable with up to $O(n)$ samples.
Next, in Section 4.2, we consider properties that are testable using $m$ samples and a circuit of size $s$, where we only require that $m$ and $s$ are subexponential in $n$. In this setting, we turn our attention from $k$-part symmetry to regularity templates, in the spirit of a landmark characterization of testable graph properties, due to [AFNS06]. In that work, it was shown that a property of dense graphs is testable from a constant number of edge queries if and only if having the property is roughly equivalent to having a regular partition from some finite set of templates. Here, a “regular partition” refers to the kind provided by the Szemerédi regularity lemma [Sze75]: a partition of the graph’s vertices into a small number of parts between which the graph’s edges are distributed pseudorandomly. By analogy, our result shows that if a property of Boolean functions is testable using $m$ samples and a circuit of size $s$, then having the property is roughly equivalent to being indistinguishable from some circuit of size $O(m^{2}s)$ in a bounded collection, where indistinguishability is defined with respect to circuits of size $s$.
Finally, in Section 4.3, we sketch a straightforward argument that our main results extend beyond the setting of testing Boolean functions from samples. Specifically, we can apply our results to distribution testing over arbitrary domains, provided that those distributions have sufficiently high min-entropy.
1.3 Technical Overview
(a) Tester $T$
(b) Simulator $\widetilde{T}_{j}$
(c) Simulator $\widetilde{T}_{j+1}$
Figure 1: Illustration of the supersimulator construction of Section 3. Plot (1(a)) depicts a deterministic 33-sample tester $T$ by the set of triples that cause it to output Accept. A one-way restriction of $T$ is shown in blue. Plot (1(b)) depicts a simulator in the sequence, with translucent regions indicating fractional estimates for $T$’s Accept region. Plot (1(c)) depicts the next simulator in the sequence, after an update has been performed based on the chosen restriction of $T$.
Our approach to proving Theorem 1.2, which states that efficient testing implies structured symmetry, will be to find a property $\mathcal{Q}$ such that $\mathcal{P}\subseteq\mathcal{Q}\subseteq\mathcal{P}_{\varepsilon}$ and membership in $\mathcal{Q}$ can be decided by a simple, structured approximation to a tester for $\mathcal{P}$. This approximation to the tester will itself be derived from simple, structured approximations to functions in $\mathcal{P}$ (or far from $\mathcal{P}$). For the latter (“simulating the oracle”) we use the complexity-theoretic regularity lemma of [TTV09], and for the former (“simulating the tester”) we use a recent enhancement of the lemma based on supersimulators [DT25].
At a high level, the complexity-theoretic regularity lemma states that any arbitrarily complex function $g:\mathcal{X}\to\{0,1\}$ has a low-complexity simulator $h:\mathcal{X}\to[0,1]$. By “simulator,” we mean that the error of $h$, namely ghg-h, is not too correlated with any function in a prespecified family $\mathcal{F}$ of distinguisher functions. Phrased differently, $g$ and $h$ are indistinguishable by $\mathcal{F}$. By “low-complexity,” we mean that $h$ is a weighted sum of a handful of functions from $\mathcal{F}$.
For example, if $\mathcal{X}=\{0,1\}^{n}$ and $\mathcal{F}$ comprises all Boolean circuits of size at most $s$, then the lemma states that any Boolean function $g:\{0,1\}^{n}\to\{0,1\}$ is indistinguishable from a circuit $h:\{0,1\}^{n}\to[0,1]$ of size [^3] We say a circuit $c$ with $n$ input bits, $m$ output bits, and $s$ logic gates computes a real-valued function $h:\{0,1\}^{n}\to[0,1]$ in binary if $h(x)=\sum_{i=1}^{m}c_{i}(x)/2^{i-1}$ for all inputs $x\in\{0,1\}^{n}$, where $c_{i}$ denotes the $i$th output bit of $c$. $O(s)$ with respect to all distinguisher circuits of size $s$. Notably, the simulator may be a constant factor larger than the distinguishers it is asked to fool. If we wish to construct simulators capable of fooling distinguishers that are larger than themselves, then we must instead use the stronger notion of supersimulators.
In what follows, we discuss the key steps of the proof in slightly more detail.
Step 1: Simulating the Oracle
First, we argue that we need not test arbitrarily complex functions $f:\{0,1\}^{n}\to\{0,1\}$ if the property of interest has a small tester $T$. Indeed, using the complexity-theoretic regularity lemma and a simple hybrid argument, we show that it suffices to test functions $\tilde{f}:\{0,1\}^{n}\to[0,1]$ that are computable by circuits slightly larger than $T$. Now, the oracle simulator returns modeled labels $\tilde{y}_{i}$ sampled from the Bernoulli distribution $\mathcal{B}(\tilde{f}(x_{i}))$, rather than real labels $y_{i}=f(x_{i})$. To the tester, however, they look no different.
Step 2: Simulating the Tester
Next, consider two $m$-sample testers $T$ and $\widetilde{T}$. Suppose that they both receive labels generated from a circuit $\tilde{f}$ that is slightly larger than both testers, and that this circuit happens to be $\{0,1\}$-valued. (Technically, $\tilde{f}$ may take fractional values, but we address this challenge in the full version of the proof.) We show that $T$ and $\widetilde{T}$ have similar probabilities of outputting Accept if $T$ and $\widetilde{T}$ are indistinguishable by a certain function derived from $\tilde{f}$, called $\tilde{f}^{\prime}$. This function receives $m$ labeled samples and checks whether they are all consistent with $\tilde{f}$:
$\tilde{f}^{\prime}(x,y)=\bm{1}\bigl[\forall i\in[m],\,y_{i}=\tilde{f}(x_{i})\bigr].$
Ideally, we would like to apply the complexity-theoretic regularity lemma once more to construct a small simulator $\widetilde{T}$ that fools $\tilde{f}^{\prime}$ for any circuit $\tilde{f}$ that is slightly larger than $T$ and $\widetilde{T}$. Indeed, if we could do this, then we would be done. The lemma would decompose $\widetilde{T}$ into a simple weighted sum of functions $\tilde{f}_{1}^{\prime},\ldots,f_{k}^{\prime}$, and taking all intersections of level sets of the low-complexity functions $\tilde{f}_{1},\ldots,\tilde{f}_{k}$ would give us the low-complexity, symmetric partition that we desire.
Step 3: Supersimulators
The discerning reader will notice the flaw in the preceding argument: we require the simulator $\widetilde{T}$ to fool distinguishers $\tilde{f}^{\prime}$ that are slightly more complex than itself, which the complexity-theoretic regularity lemma cannot provide! To overcome this obstacle, we apply supersimulators, which are a powerful strengthening of the lemma that constructs simulators capable of fooling significantly more complex distinguishers.
At a high level, the supersimulator construction we employ is iterative, generating a sequence of approximations $\widetilde{T}_{1},\widetilde{T}_{2},\ldots$ to the original tester $T$. At the $j$th step, we derive $\widetilde{T}_{j+1}$ from $\widetilde{T}_{j}$ by first choosing a function $\tilde{f}_{j}$ slightly larger than $\widetilde{T}_{j}$ and then adding an appropriate multiple of $\tilde{f}_{j}^{\prime}$ to $\widetilde{T}_{j}$. Specifically, the function $\tilde{f}_{j}$ will a combination of a handful of one-way restrictions of $\widetilde{T}_{j}$, which are obtained by hard-wiring all but one input to the current simulator $\widetilde{T}_{j}$. We illustrate the construction in Figure 1. We discuss supersimulators further in Section 2, give a full proof of Theorems 1.2 and 1.3 in Section 3, and prove our various extensions in Section 4.
1.4 Related Work
Our paper extends a recent line of work investigating the interplay between classical results in pseudorandomness (specifically, the structure-vs-randomness paradigm) and recent results in the algorithmic fairness literature. In this work, we do so through the lens of property testing.
Structure vs Randomness
There is a large body of work devoted to decomposing complex objects into their structured and pseudorandom components. An early result in this space is Szemerédi’s regularity lemma [Sze75]. The result roughly states that any large, dense graph can be split into a (rather large) handful of parts between which the graph’s edges are distributed pseudorandomly. There is a vast literature on the regularity lemma and its variants, producing breakthrough results in pseudorandomness and additive combinatorics to this day [KM23, JLL+25]. For more, see the surveys [Tao07, Zha23].
An especially relevant result in this field is the complexity-theoretic regularity lemma of [TTV09], which gave a unified perspective on such disparate topics as the Frieze-Kannan weak regularity lemma for graphs [FK96, FK99], Impagliazzo’s hardcore lemma [Imp95], and the dense model theorem [GT08, TZ08, RTTV08]. It also led to a deeper understanding of computational entropy [VZ12, VZ13, Zhe14] and techniques for leakage simulation in cryptography [JP14].
Algorithmic Fairness
In the context of algorithmic fairness for machine learning systems, modern concepts like outcome indistinguishability [DKR+21] and multicalibration [HKRR18, KNRW18], studied further in work on omniprediction [GKR+22, GHK+23], can be viewed as stronger, constructive versions of complexity-theoretic regularity that utilize practical learning-theoretic primitives.
The connection between these modern tools and older notions of regularity was made explicit by [DLLT23, CDV24], and this point of view has proven to be fruitful. Indeed, these works and their sequelae have led to new insights into graph regularity, hardcore set construction, dense models, omniprediction, computational hardness and entropy, and the computational indistinguishability of product distributions [MPV25, CGKR25, HV25, DT25]. The present work similarly builds on the ideas from this line of work, but now with a view toward property testing. In particular, supersimulation [DT25], which is our main technical tool, is closely related to both outcome indistinguishability and multicalibration.
Property Testing
The study of property testing writ large was initiated by [RS96, GGR96]. Within this field, several works have attempted to characterize which properties are testable under various constraints on the tester’s power. For example, in the graph context, [AFNS06] proved, roughly speaking, that a property of dense graphs is testable if and only if it can be determined from a Szemerédi regular partition of the graph. In some sense, this result was a capstone to a large body of work devoted to understanding the testability of graph properties, including monotone properties [AS05b], hereditary properties [AS05a], and more.
In the present paper, we do not work with dense graphs, but instead test properties of a certain class of dense distributions defined by Boolean functions. As already discussed, in this context, the closest related work to ours is the characterization of constant-sample testability in terms of constant-part symmetry, due to [BY19]. Even earlier, [KS08, Sud10] pioneered the idea that the symmetries of a property play a central role in understanding its testability.
Our work concerns the computational complexity of property testing, which has recently received renewed interest [FPR26]. For the related but simpler task of efficiently determining which of two distributions $\mathcal{D}_{0}$ vs $\mathcal{D}_{1}$ generated an observed set of $m$ samples, the recent work of [MPV25] used multicalibration to argue that densities on a low-complexity partition of the domain contain all the relevant information (see also [DT25]). Our results have a similar flavor, but we test for membership in sets $\mathcal{P}$ vs $\neg\mathcal{P}_{\varepsilon}$ that may contain doubly exponentially many distributions, rather than just one each.
2 Preliminaries
In this paper, $\mathcal{X}$ denotes an arbitrary finite set, $\Delta(\mathcal{X})$ denotes the set of probability distributions on $\mathcal{X}$, and $\{\mathcal{X}\to\mathcal{Y}\}$ denotes the set of functions from $\mathcal{X}$ to a set $\mathcal{Y}$. Given $\mathcal{F}\subseteq\{\mathcal{X}\to\mathbb{R}\}$ and $c\in\mathbb{R}$, let $c\cdot\mathcal{F}$ denote the set of functions $c\cdot f$ for $f\in\mathcal{F}$. Let $\mathcal{F}=(-1)\cdot\mathcal{F}$ and let $\pm\mathcal{F}=\mathcal{F}\cup$-$\mathcal{F}$. Let $\mathcal{B}(p)$ denote the Bernoulli distribution with parameter $p\in[0,1]$. Finally, let $t]_{a}^{b}$ denote the projection of $t\in\mathbb{R}$ onto the interval $[a,b]$.
Property Testing
A property $\mathcal{P}$ is a set of Boolean functions $f:\{0,1\}^{n}\to\{0,1\}$. We say that $f$ has the property $\mathcal{P}$ if $f\in\mathcal{P}$. The distance between two functions is the fraction of inputs on which they disagree. We write $f\in\mathcal{P}_{\varepsilon}$ if $f$ is $\varepsilon$-close to some $g\in\mathcal{P}$. While property testing can be studied in either a query-based or sample-based access model, we focus exclusively on the latter perspective, which was introduced by [GGR96]:
Definition 2.1 (Sample-Testable Property).
Let $\mathcal{P}$ be a property of Boolean functions. We say that $\mathcal{P}$ is sample-testable with proximity parameter $\varepsilon>0$ if there is a randomized circuit $T$ of size $s$ that receives as input $m$ independent samples $x_{i}\sim\{0,1\}^{n}$ and their labels $y_{i}=f(x_{i})$, always outputs either Accept or Reject, and meets the following two requirements:
  • If $f$ has the property $\mathcal{P}$, then $T$ outputs Accept with probability at least $2/3$.
  • If $f$ is $\varepsilon$-far from having $\mathcal{P}$, then $T$ outputs Reject with probability at least $2/3$.
In both conditions, the probability is computed over randomness in the sample and internal to $T$.
More formally, we will model the tester as a deterministic function $T:(\mathcal{X}\times\{0,1\})^{m}\times\{0,1\}^{\ell}\to\{0,1\}$ that receives as input $m$ labeled samples $(x_{i},y_{i})\in\mathcal{X}\times\{0,1\}$ and a uniform $\ell$-bit random seed $r\in\{0,1\}^{\ell}$. Often, we will write $x\in\mathcal{X}^{m}$ and $y\in\{0,1\}^{m}$ and use the abbreviation
$T(x,y,r)=T\bigl((x_{1},y_{1}),\ldots,(x_{m},y_{m}),r\bigr).$
We say that $T$ “accepts” when it outputs $1$, and “rejects” when it outputs 0. Sometimes, it will be convenient to work directly with the expected value of $T$ over its internal randomness, or, equivalently, the mean function $\bar{T}:(\mathcal{X}\times\{0,1\})^{m}\to[0,1]$ defined by the formula
$\bar{T}(x,y)=2^{-\ell}\sum_{r\in\{0,1\}^{\ell}}T(x,y,r).$
When $\mathcal{X}=\{0,1\}^{n}$, we will often discuss the circuit size of the tester, by which we mean $T$, not $\bar{T}$.
Structured Symmetry
A property $\mathcal{P}$ of Boolean functions is $k$-part symmetric if there is a partition of $\{0,1\}^{n}$ into disjoint parts $S_{1},\ldots,S_{k}$ such that $\mathcal{P}$ is invariant under permutations of the domain within each part. Equivalently, whether or not a function $f$ has the property $\mathcal{P}$ can be completely determined from the $k$ scalar densities $\operatorname*{\mathbb{E}}[f(x)\,|\,x\in S_{j}]$, where the expectation is computed over a random input $x\sim\{0,1\}^{n}$. In general, such parts need not have any special structure, and may be very complex. In contrast, we say that a part $S_{j}$ has size complexity at most $s$ if there is a circuit of size $s$ that decides whether or not a given input belongs to $S_{j}$. We say that a partition $\mathcal{P}$ has size complexity at most $s$ if there is a circuit of size at most $s$ computing its classification function, which computes the index of the part $S_{j}$ containing a given input $x$.
Supersimulators
Our key technical tool is the construction of supersimulators [DT25], which strengthens the complexity-theoretic regularity lemma [TTV09] by designing regular simulators that fool distinguishers more powerful than themselves. First, we define indistinguishability formally.
Definition 2.2 (Regularity and Indistinguishability).
Given a family $\mathcal{F}\subseteq\{\mathcal{X}\to[0,1]\}$, an error parameter $\delta>0$, and two functions $g,h:\mathcal{X}\to[0,1]$, we say that $h$ is a $(\mathcal{F},\delta)$-regular simulator for $g$ under $\mathcal{D}$ if for all distinguisher functions $f\in\mathcal{F}$,
[…]
Equivalently, we say that $g$ and $h$ are $(\mathcal{F},\delta)$-indistinguishable (when clear from context, we will omit the phrase “under $\mathcal{D}$”). When $\mathcal{D}$ is the uniform distribution over $\mathcal{X}=\{0,1\}^{n}$ and $\mathcal{F}$ is the collection of all Boolean circuits of size at most $s$, we say that $h$ is an $(s,\delta)$-regular simulator for $g$, or that the functions are $(s,\delta)$-indistinguishable.
Given any target function $g:\mathcal{X}\to[0,1]$, there exists a trivial $(\mathcal{F},\delta)$-regular simulator for $g$, namely $h=g$. The complexity-theoretic regularity lemma guarantees the existence of a much better simulator, whose complexity does not scale with $\mathcal{X}$ or $g$, but rather depends only on $\mathcal{F}$ and $\delta$. We will state the lemma in terms of the distinguisher family’s structured sums $\mathcal{S}_{k,\delta}(\mathcal{F})$.
Definition 2.3 (Structured Sums).
If $\mathcal{F}\subseteq\{\mathcal{X}\to[-1,1]\}$, then $\mathcal{S}_{k,\delta}(\mathcal{F})$ is the set of functions
$h(x)=\Bigl[\delta\cdot\bigl(f_{1}(x)+\cdots+f_{k}(x)\bigr)\Bigr]_{0}^{1}$
for some $f_{1},\ldots,f_{k}\in\pm\mathcal{F}$. (Recall that $\cdot]_{0}^{1}$ projects onto $[0,1]$.) Let $\mathcal{S}_{<k,\delta}(\mathcal{F})=\bigcup_{j<k}\mathcal{S}_{j,\delta}(\mathcal{F})$.
Lemma 2.4 (Complexity-Theoretic Regularity [TTV09]).
Fix $\mathcal{D}\in\Delta(\mathcal{X})$, $\mathcal{F}\subseteq\{\mathcal{X}\to[-1,1]\}$, and $\delta>0$. Every $g:\mathcal{X}\to[0,1]$ has an $(\mathcal{F},\delta)$-regular simulator $h\in\mathcal{S}_{<(2/\delta^{2}),(\delta/2)}(\mathcal{F})$.
Lemma 2.4 may produce a simulator $h$ that is more complex than the functions in $\mathcal{F}$, but for our purposes, we will need $h$ to be a supersimulator, which fools distinguishers more complex than itself. We formalize this in terms of a growth function $\mathcal{G}$ that takes as input any function $h:\mathcal{X}\to[0,1]$ and outputs the distinguisher family $\mathcal{G}(h)\subseteq\{\mathcal{X}\to[-1,1]\}$ that we would like $h$ to fool. Now, the definition of structured sums must be adjusted accordingly.
Definition 2.5.
If $\mathcal{G}$ is a growth function, then $\mathcal{S}_{k,\delta}(\mathcal{G})$ is the set of functions
$h_{k}(x)=\Bigl[\delta\cdot\bigl(f_{1}(x)+\cdots+f_{k}(x)\bigr)\Bigr]_{0}^{1}$
where each $f_{j}$ belongs to the previous family $\pm\mathcal{G}(h_{j-1})$, and $h_{0}=0$. Let $\mathcal{S}_{<k,\delta}(\mathcal{G})=\bigcup_{j<k}\mathcal{S}_{j,\delta}(\mathcal{G})$.
Lemma 2.6 (Supersimulators [DT25]).
Fix $\mathcal{D}\in\Delta(\mathcal{X})$, a growth function $\mathcal{G}$, and $\delta>0$. Every $g:\mathcal{X}\to[0,1]$ has a $(\mathcal{G}(h),\delta)$-regular simulator $h\in\mathcal{S}_{<(2/\delta^{2}),(\delta/2)}(\mathcal{G})$.
Note that Lemma 2.4 is a special case of Lemma 2.6 corresponding to a constant growth function that always outputs $\mathcal{F}$. We remark that the statement of Lemma 2.6 differs slightly from that of [DT25] because Lemma 2.6 uses a single projection onto the interval $[0,1]$, as opposed to a nested sequence of projections. For completeness, we provide a short proof of this version in Appendix A. In fact, our argument also yields a slightly simpler proof of Lemma 2.4 than in [TTV09], which required some careful case analysis to handle the projection. Simple proofs of the complexity-theoretic regularity lemma without this case analysis have already appeared (e.g. implicitly in [HKRR18], as observed by [CDV24]), but these versions also require a sequence of nested projections, unlike Lemma 2.4, which more closely resembles the original version from [TTV09].
3 Finding Structured Symmetry
In this section, we prove Theorems 1.3 and 1.2, which comprise our main equivalence between efficient testability and structured symmetry. As discussed in the introduction, this result strengthens the main theorem of [BY19] by capturing the relationship between the tester’s complexity and the structure of the property’s invariant sets. For convenience, we recall the statements of these two theorems below. The first, Theorem 1.2, is the harder of the two, showing that efficient testing implies structured symmetry.
See 1.2
Like the main theorem of [BY19], our Theorem 1.2 is most interesting for properties testable from at most $O(\log n)$ samples, where $2^{n}$ is the domain size. We will extend our results to properties testable with larger sample sizes in subsequent sections. The second theorem, Theorem 1.3, establishes the converse to Theorem 1.2 by showing that structured symmetry implies efficient testing.
See 1.3
To see why Theorem 1.3 is indeed a converse to Theorem 1.2, note that if $\mathcal{P}\subseteq\mathcal{Q}\subseteq\mathcal{P}_{\varepsilon_{1}}$ for some $\varepsilon_{1}>0$, then $\mathcal{Q}_{\varepsilon_{2}}\subseteq\mathcal{P}_{\varepsilon_{1}+\varepsilon_{2}}$ for any $\varepsilon_{2}>0$. Consequently, any tester for $\mathcal{Q}$ with proximity $\varepsilon_{2}$ is also a tester for $\mathcal{P}$ with proximity $\varepsilon_{1}+\varepsilon_{2}$. Thus, when combined, Theorems 1.2 and 1.3 imply that a property of Boolean functions is constant-sample testable with a small circuit if and only if it is close to having constant-part structured symmetry.
While the proof of Theorem 1.3 is quite simple, the proof of Theorem 1.2 is not, requiring two key lemmas. We call them the oracle simulation lemma and the tester simulation lemma, and we prove them in Sections 3.1 and 3.2, respectively. While both lemmas involve regular simulators, they differ in terms of which objects play the roles of the target function and distinguisher family. For the oracle simulation lemma, the target function is the one we wish to test for $\mathcal{P}$, which defines the example oracle, and the distinguishers are derived from the tester. For the tester simulation lemma, the tester plays the role of the target function, and the distinguisher family comprises low-complexity approximations to the functions to be tested for $\mathcal{P}$. In Section 3.3, we combine these lemmas to prove Theorem 1.2, and we also prove Theorem 1.3.
3.1 Simulating the Oracle
In this section, we show that the probability that $T$ accepts or rejects any particular example oracle, which is defined by a distribution $\mathcal{D}\in\Delta(\mathcal{X})$ and a target function $f:\mathcal{X}\to[0,1]$, remains roughly the same upon replacing $f$ with any $(\mathcal{R}(T),\delta)$-regular simulator $\tilde{f}$, where $\mathcal{R}(T)$ is the simple distinguisher family described in Definition 3.1 below. Specifically, $\mathcal{R}(T)$ is a family of one-way restrictions of $T$, obtained by hard-wiring fixed values for all but one of its inputs. Later, when we take $T$ to be a size-$s$ circuit, it will be clear that every distinguisher of this form is also a size-$s$ circuit.
The main result of this section, which we state formally in Lemma 3.2, is significant because of the way it facilitates the construction of additional property testers. To illustrate this informally, suppose that a property is testable by a simple circuit $T$ and that we would like to argue that some other function $\widetilde{T}$ is also a valid tester. One way to do this would be to argue that $\widetilde{T}$ behaves similarly to $T$ for all possible example oracles, but this may be difficult. Lemma 3.2 shows it suffices to check that $\widetilde{T}$ behaves similarly to $T$ for example oracles defined by low-complexity target functions $\tilde{f}$. We note, however, that the result of this section will be stated without the language of property testing, as it depends only on the function $T$.
Before proceeding with the statement and proof, we first define the relevant distinguisher class. The definition is precisely the one needed to enable a certain hyrbid argument that we plan to carry out. We state the definition in a a general manner so that it is applicable to both the actual tester $T:(\mathcal{X}\times\{0,1\})^{m}\times\{0,1\}^{\ell}\to\{0,1\}$ and its mean function $\bar{T}:(\mathcal{X}\times\{0,1\})^{m}\to[0,1]$, which averages over the choice of the random seed $r\in\{0,1\}^{\ell}$.
Definition 3.1 (Restriction Distinguishers).
Given a function $T:(\mathcal{X}\times\{0,1\})^{m}\times\{0,1\}^{\ell}\to[0,1]$, consider the function $T_{x_{\neq i},y,r}:\mathcal{X}\to[0,1]$ which hard-wires all inputs to $T$ except for $x_{i}$:
$T_{x_{\neq i},y,r}(x)=T\bigl((x_{1},y_{1}),\ldots,(x_{i-1},y_{i-1}),(x,y_{i}),(x_{i+1},y_{i+1}),\ldots,(x_{m},y_{m}),r\bigr).$
We define $\mathcal{R}(T)$ to be the set of these one-way restriction functions $T_{x_{\neq i},y,r}$ for all indices $i\in[m]$, sequences $x_{\neq i}$ comprising values $x_{j}\in\mathcal{X}$ for each $j\neq i$, labels $y\in\{0,1\}^{m}$, and seeds $r\in\{0,1\}^{\ell}$.
In the special case that $T$ is Boolean-valued, each function $T_{x_{\neq i},y,r}$ has codomain $\{0,1\}$. With this definition in hand, we are now ready to state the main result of this section:
Lemma 3.2 (Oracle Simulation).
Fix a function $T:(\mathcal{X}\times\{0,1\})^{m}\times\{0,1\}^{\ell}\to[0,1]$, a distribution $\mathcal{D}\in\Delta(\mathcal{X})$, and $\delta>0$. If $\tilde{f}$ is an $(\mathcal{R}(T),\delta)$-regular simulator for $f:\mathcal{X}\to[0,1]$, then
$\Bigl\lvert\operatorname*{\mathbb{E}}\bigl[T(x,y,r)\bigr]-\operatorname*{\mathbb{E}}\bigl[T(x,\tilde{y},r)\bigr]\Bigr\rvert\leq 2m\delta,$
where $x_{i}\sim\mathcal{D}$ and $y_{i}|x_{i}\sim\mathcal{B}(f(x_{i}))$ and $\tilde{y}_{i}|x_{i}\sim\mathcal{B}(\tilde{f}(x_{i}))$ for each $i\in[m]$ and $r\sim\{0,1\}^{\ell}$.
A remark is in order regarding the random variables $y_{i}$ and $\tilde{y}_{i}$. In the language of Outcome Indistinguishability (OI) [DKR+21] from the algorithmic fairness literature, $y_{i}$ and $\tilde{y}_{i}$ correspond precisely to real and modeled outcomes of an individual represented by the features $x_{i}$. From this point of view, $f$ defines the ground-truth or Bayes optimal conditional probability distribution, and $\tilde{f}$ corresponds to a predictor of $f$ satisfying no-access OI with respect to $\mathcal{R}(T)$.
The proof of Lemma 3.2 involves two simple components. The first component is a hybrid argument, similar to one used in recent work on multicalibration-based characterizations of the indistinguishability of product distributions [MPV25, DT25]. The other component is a standard transformation between distinguishers that receive a labeled or unlabeled input.
Proof.
First, for each integer $1\leq i\leq m$, define
$T_{i}(x,y)=T\bigl((x_{1},y_{1}),\ldots,(x_{i-1},y_{i-1}),(x,y),(x_{i+1},\tilde{y}_{i+1}),\ldots,(x_{m},\tilde{y}_{m}),r\bigr).$
Observe that the function $T_{i}:\mathcal{X}\times\{0,1\}\to\{0,1\}$ depends on the values of $x_{j}$ for all $j\neq i$, the values of $y_{j}$ for all $j<i$, the values of $\tilde{y}_{j}$ for all $j>i$, and the tester’s random seed $r$. Next, for each $0\leq i\leq m$, let the (deterministic) scalar $a_{i}$ denote the expected output of the tester $T$ under the $i$th hybrid distribution, in which the first $i$ labels are real and the remaining mim-i labels are modeled. More formally, for $1\leq i\leq m-1$, we define
$a_{i}=\operatorname*{\mathbb{E}}\bigl[T_{i}(x_{i},y_{i})\bigr]=\operatorname*{\mathbb{E}}\bigl[T_{i+1}(x_{i+1},\tilde{y}_{i+1})\bigr].$
When $i=m$, we define $a_{m}$ via the first of these two expressions, and when $i=0$, we define $a_{0}$ via the second of these two expressions. Ultimately, our goal is to bound $\lvert a_{m}-a_{0}\rvert\leq 2m\delta$. By the triangle inequality, it suffices to show that $\lvert a_{i}-a_{i-1}\rvert\leq 2\delta$ for each index $i\in[m]$. For this, we condition on $x_{j}$ for all $j\neq i$ and $y_{j}$ for all $j<i$ and $\tilde{y}_{j}$ for all $j>i$ (i.e. everything except for $x_{i}$, $y_{i}$ and $\tilde{y}_{i}$):
[…]
Note that the inner expectation is only over the randomness in the $i$th coordinate (i.e. $x_{i}$, $y_{i}$, and $\tilde{y}_{i}$), since we have conditioned on everything else. The outer expectation is over the randomness of these other variables, namely $x_{\neq i}$, $y_{<i}$, $\tilde{y}_{>i}$ and $r$.
At this point, we have computed an upper bound on $\lvert a_{i}-a_{i-1}\rvert$ in terms of the distinguishing advantage of a circuit $T_{i}$ derived from hard-wiring all but two of the inputs to $T$. However, we are not quite done until we have hard-wired all but one input, since this is the form required by $\mathcal{R}(T)$. To this end, we apply a standard transformation to $T_{i}$. First, since $y_{i}|x_{i}\sim\mathcal{B}(f(x_{i}))$, we rewrite the conditional expectation of $T_{i}(x_{i},y_{i})$ in terms of $f(x_{i})$ instead of $y_{i}$:
$\operatorname*{\mathbb{E}}\bigl[T_{i}(x_{i},y_{i})\bigr]=\operatorname*{\mathbb{E}}\bigl[T_{i}(x_{i},0)+(T_{i}(x_{i},1)-T_{i}(x_{i},0))f(x_{i})\bigr].$
Since $\tilde{y}_{i}|x_{i}\sim\mathcal{B}(\tilde{f}(x_{i}))$, a similar formula holds for $\tilde{y}_{i}$ with $\tilde{f}$ in place of $f$:
$\operatorname*{\mathbb{E}}\bigl[T_{i}(x_{i},\tilde{y}_{i})\bigr]=\operatorname*{\mathbb{E}}\bigl[T_{i}(x_{i},0)+(T_{i}(x_{i},1)-T_{i}(x_{i},0))\tilde{f}(x_{i})\bigr].$
We observe that these two formulas remain true even if we condition on the values of $x_{i}$, $x_{\neq i}$, $y_{<i}$, $\tilde{y}_{>i}$, and $r$. Therefore, subtracting the two equations yields
[…]
The functions $T_{i}(x,1)$ and $T_{i}(x,0)$ clearly belong to $\mathcal{R}(T)$. Since $\tilde{f}$ is an $(\mathcal{R}(T),\delta)$-regular simulator for $f$, we have $\lvert a_{i}-a_{i-1}\rvert\leq 2\delta$, so we conclude that $\lvert a_{m}-a_{0}\rvert\leq 2m\delta$. ∎
3.2 Simulating the Tester
In this section, we show that replacing $T$ with a suitable supersimulator $\widetilde{T}$ only slightly affects our chance of accepting or rejecting any sufficiently low-complexity example oracle. This promise problem, in which we assume that the example oracle has low complexity, is a natural variant of the property testing framework defined in Section 2. It is motivated by our main result from the preceding section, which showed that for some purposes, it suffices to restrict attention to oracles defined by low-complexity target functions. Later, in Section 3.3, we will show how to use $\widetilde{T}$ to establish the structured symmetry of a property under consideration.
The main result of this section, which we state formally in Lemma 3.4, will require the notion of consistency indicators of a distinguisher family.
Definition 3.3 (Consistency Indicators).
Given a function family $\mathcal{F}\subseteq\{\mathcal{X}\to[0,1]\}$, consider the family of consistency indicators $\Gamma_{m}(\mathcal{F})\subseteq\{(\mathcal{X}\times\{0,1\})^{m}\to\{0,1\}\}$, which take as input $m$ labeled pairs $(x_{i},y_{i})$ and checks whether they are all consistent with some function in $\mathcal{F}$, after thresholding:
$\Gamma_{m}(\mathcal{F})=\Bigl\{(x,y)\mapsto\bm{1}\bigl[\forall i\in[m],\,y_{i}=\bm{1}[f(x_{i})\geq t_{i}]\bigr]\,\Big|\,f\in\mathcal{F},\,t_{1},\ldots,t_{m}\in\mathbb{R}\Bigr\}.$
As usual, $(x,y)$ is our abbreviation for the $m$-tuple of pairs $(x_{i},y_{i})$.
Much like our proof of the oracle simulation lemma in Section 3.1, our proof of Lemma 3.4 is based on indistinguishability. While we will eventually take the simulator in Section 3.1 to be the one provided by the complexity-theoretic regularity lemma, the simulator in the present section will eventually come from the supersimulators lemma. Interestingly, in Section 3.1, the oracle played the role of the object to be simulated, and the property tester played the role of the distinguisher. In contrast, in this section, the tester shall play the role of the object to be simulated, and the oracle shall play the role of the distinguisher.
Lemma 3.4 (Tester Simulation).
Fix a distribution $\mathcal{D}\in\Delta(\mathcal{X})$ and a family $\mathcal{F}\subseteq\{\mathcal{X}\to[0,1]\}$. If $\widetilde{T}$ is a $(\Gamma_{m}(\mathcal{F}),\gamma)$-regular simulator for $\bar{T}:(\mathcal{X}\times\{0,1\})^{m}\to[0,1]$, then for all $\tilde{f}\in\mathcal{F}$,
$\Bigl\lvert\operatorname*{\mathbb{E}}\bigl[\bar{T}(x,\tilde{y})\bigr]-\operatorname*{\mathbb{E}}\bigl[\widetilde{T}(x,\tilde{y})\bigr]\Bigr\rvert\leq 2^{m}\gamma,$
where $x_{1},\ldots,x_{m}\overset{\textit{iid}}{\sim}\mathcal{D}$ and $\tilde{y}_{i}|x_{i}\sim\mathcal{B}(\tilde{f}(x_{i}))$.
Lemma 3.4 can be viewed as a generalization of a technical result from [BY19]. While our result applies to a general family $\mathcal{F}\subseteq\{\mathcal{X}\to[0,1]\}$, the technical result from [BY19] corresponds to the special case of the class $\mathcal{F}_{0}=\{\mathcal{X}\to\{0,1\}\}$. Note that $\mathcal{F}$ may contain $[0,1]$-valued functions, which is why we need to incorporate the notion of consistency indicators $\Gamma_{m}$ after thresholding. These considerations were not needed for the case of $\mathcal{F}_{0}$, which contains only Boolean functions.
While [BY19] obtain their regular approximation to $T$ using a hypergraph regularity lemma, we will eventually acquire $\widetilde{T}$ from the supersimulators lemma, thus circumventing the hypergraph-based formalism.
Proof of Lemma 3.4.
We first rewrite the labels $\tilde{y}_{i}$ in a convenient form. Specifically, we write
$\tilde{y}_{i}=\bm{1}[\tilde{f}(x_{i})\geq t_{i}]$
[…]
$\bar{T}(x,\tilde{y})=2^{m}\cdot\operatorname*{\mathbb{E}}_{z}\bigl[\bar{T}(x,z)\,\bm{1}[\tilde{y}=z]\,\big|\,x,t\bigr].$ (*)
Next, observe that the indicator function $\bm{1}[\tilde{y}=z]$ can be written as $\bm{1}[z=\bm{1}[\tilde{f}(x)\geq t]]$. (Here, the condition $\tilde{f}(x)\geq t$ should be read coordinate-wise.) When viewed as a function of $x$ and $z$, this function belongs to the distinguisher family $\Gamma_{m}(\mathcal{F})$, by definition. Note that this is a random function, which depends on the thresholds $t$. By assumption, $\widetilde{T}$ is a $(\Gamma_{m}(\mathcal{F}),\gamma)$-regular simulator for $\bar{T}$. Therefore, taking the expectation over $x$,
[…]
Finally, taking the expectation over $t$ and using the identity ()(*) yields
$\Bigl\lvert\operatorname*{\mathbb{E}}_{x,t}\bigl[\bar{T}(x,\tilde{y})\bigr]-\operatorname*{\mathbb{E}}_{x,t}\bigl[\widetilde{T}(x,\tilde{y})\bigr]\Bigr\rvert\leq 2^{m}\gamma.\qed$
3.3 Constructing the Partition
In this section, we prove Theorems 1.2 and 1.3. First, we instantiate Lemma 3.2 (oracle simulation) with Lemma 2.4 (complexity-theoretic regularity). Next, we instantiate Lemma 3.4 (tester simulation) with Lemma 2.6 (supersimulators). Finally, we turn our attention to small circuit testers and use the fact that operations like restrictions and thresholding roughly preserve circuit size.
See 1.2
Proof of Theorem 1.2.
Let $T:(\mathcal{X}\times\{0,1\})^{m}\times\{0,1\}^{\ell}\to\{0,1\}$ be a valid tester for $\mathcal{P}$. This means that $T$ outputs Accept with probability at least $2/3$ when $f\in\mathcal{P}$ and outputs Reject with probability at least $2/3$ when $f$ is $\varepsilon$-far from $\mathcal{P}$. As usual, these probabilities are with respect to independent samples $x_{i}\sim\mathcal{D}$ with labels $y_{i}=f(x_{i})$, and a uniform $\ell$-bit random seed $r\in\{0,1\}^{\ell}$. As usual, let $\bar{T}(x,y)$ denote the expected value of $T(x,y,r)$ over $r$. We proceed in several steps.
Step 1: Constructing the Supersimulator
In order to approximate $\mathcal{P}$ by a property $\mathcal{Q}$ with structured symmetry, we will first construct a supersimulator $\widetilde{T}$ for $T$, and then extract the desired partition from the inner structure of $\widetilde{T}$. To this end, fix $\delta,\gamma>0$. Let $\mathcal{G}$ be the growth function that takes as input a function $T^{\prime}:(\mathcal{X}\times\{0,1\})^{m}\to[0,1]$ and outputs the distinguisher family
$\mathcal{G}(T^{\prime})=\Gamma_{m}\Bigl(\mathcal{S}_{<(2/\delta^{2}),(\delta/2)}\bigl(\mathcal{R}(T)\cup\mathcal{R}(T^{\prime})\bigr)\Bigr).$
By Lemma 2.6 (supersimulators), $\bar{T}$ has a $(\mathcal{G}(\widetilde{T}),\gamma)$-regular supersimulator
$\widetilde{T}\in\mathcal{S}_{<(2/\gamma^{2}),(\gamma/2)}(\mathcal{G}).$
By Lemma 2.4 (complexity-theoretic regularity), $f$ has an $(\mathcal{R}(T)\cup\mathcal{R}(\widetilde{T}),\delta)$-regular simulator
[…]
As usual, we let $\tilde{y}_{i}$ denote the modeled labels generated from $\tilde{f}$, which means that $\tilde{y}_{i}|x_{i}\sim\mathcal{B}(\tilde{f}(x_{i}))$.
Step 2: Applying the Two Key Lemmas
Next, we will use Lemma 3.2 (oracle simulation) and Lemma 3.4 (tester simulation) to show that $T$ and $\widetilde{T}$ have similar probabilities of outputting Accept regardless of the labeling function $f:\mathcal{X}\to\{0,1\}$. To make the argument more concise, we will write $a\approx_{\rho}b$ if two scalars $a,b\in\mathbb{R}$ differ in absolute value by at most $\rho$. First, since $\tilde{f}$ is a $(\mathcal{R}(T),\delta)$-regular simulator for $f$, Lemma 3.2 implies
[…]
Next, since $\widetilde{T}$ is a $(\mathcal{G}(\widetilde{T}),\gamma)$-regular simulator for $\bar{T}$, Lemma 3.4 implies
[…]
Finally, since $\tilde{f}$ is a $(\mathcal{R}(\widetilde{T}),\delta)$-regular simulator for $f$, applying Lemma 3.2 again yields
$\operatorname*{\mathbb{E}}\bigl[\widetilde{T}(x,\tilde{y})\bigr]\approx_{2m\delta}\operatorname*{\mathbb{E}}\bigl[\widetilde{T}(x,y)\bigr].$
Combining these three steps, we deduce that
$\operatorname*{\mathbb{E}}\bigl[T(x,y,r)\bigr]\approx_{4m\delta+2^{m}\gamma}\operatorname*{\mathbb{E}}\bigl[\widetilde{T}(x,y)\bigr].$
In other words, the probability that $T$ outputs Accept differs from the expected output of $\widetilde{T}$ by at most $4m\delta+2^{m}\gamma$, which is less than $1/6$ for appropriately chosen $\delta=\Theta(1/m)$ and $\gamma=\Theta(1/2^{m})$.
Step 3: Defining the New Property
Having related $\widetilde{T}$ to $T$, we define the property
$\mathcal{Q}=\biggl\{f:\mathcal{X}\to\{0,1\}\;\bigg|\;\operatorname*{\mathbb{E}}\bigl[\widetilde{T}(x,y)\bigr]\geq\frac{1}{2}\biggr\},$
where, as usual, the expectation is over $m$ independent samples $x_{i}\sim\mathcal{D}$, where $y_{i}=f(x_{i})$. In other words, we say that a function $f$ has the property $\mathcal{Q}$ if the expected output of the simulated tester $\widetilde{T}$ is at least $1/2$ on samples labeled by $f$.
We claim that $\mathcal{P}\subseteq\mathcal{Q}\subseteq\mathcal{P}_{\varepsilon}$. Indeed, this follows immediately from the previously established relationship between $T$ and $\widetilde{T}$. In slightly more detail, suppose that $f\in\mathcal{P}$. Then, since $T$ is a valid tester for $\mathcal{P}$, it outputs Accept with probability at least $2/3$. Consequently, the expected output of $\widetilde{T}$ is at least 2/31/6=1/22/3-1/6=1/2. Thus, $f\in\mathcal{Q}$. A similar argument shows that $\mathcal{Q}\subseteq\mathcal{P}_{\varepsilon}$. At this point, we have shown that $\mathcal{P}\subseteq\mathcal{Q}\subseteq\mathcal{P}_{\varepsilon}$, where membership in $\mathcal{Q}$ depends only on the probability of acceptance by $\widetilde{T}$.
Step 4: Defining the Partition
Finally, we show that $\mathcal{Q}$ exhibits the desired structured symmetry. For this, recall that the supersimulator lemma states that $\widetilde{T}$ has the form
$\widetilde{T}=\widetilde{T}_{k}=\biggl[\frac{\gamma}{2}\bigl(F_{1}+\cdots+F_{k}\bigr)\biggr]_{0}^{1}$
for some $k<2/\gamma^{2}$, where each term $F_{j}$ belongs to $\pm\mathcal{G}(\widetilde{T}_{j-1})$. In other words, $F_{j}$ is a (possibly negated) consistency indicator for some function
$f_{j}\in\mathcal{S}_{<(2/\delta^{2}),(\delta/2)}\bigl(\mathcal{R}(T)\cup\mathcal{R}(\widetilde{T}_{j-1})\bigr).$
Consequently, the expected output of $\widetilde{T}$ depends only on the density of the labeling function on the following collection of sets, for various indices $i\in[m]$ and $j\in[k]$ and fixed thresholds $t_{ij}\in[0,1]$:
$S_{ij}=\{x\in\mathcal{X}\mid f_{j}(x)\geq t_{ij}\}.$
Taking all $2^{mk}\leq 2^{2^{O(m)}}$ intersections of the sets $S_{ij}$ yields a partition witnessing the $2^{2^{O(m)}}$-part symmetry of $\mathcal{Q}$.
Finally, we address the partition complexity of $\mathcal{Q}$. For this, let $\mathcal{D}$ be the uniform distribution on $\mathcal{X}=\{0,1\}^{n}$ and suppose that $T$ is computable by a circuit of size $s$. In this case, the restrictions $\mathcal{R}(T)$ obtained by hard-wiring various inputs to $T$ are also computable by circuits of size $s$. Therefore, the subsequent Lemma 3.5 (with our choices of $\delta=\Omega(1/m)$, $\gamma=\Omega(1/2^{m})$, and $k=O(1/\gamma^{2})$) shows that $\mathcal{Q}$ has partition complexity at most $2^{O(m)}s$, concluding the proof of Theorem 1.2. ∎
Lemma 3.5 (Counting Circuit Gates).
Fix $\delta,\gamma>0$, functions $f_{j}:\mathcal{X}\to[0,1]$, thresholds $t_{ij}\in[0,1]$, and signs $\sigma_{j}\in\{-1,+1\}$ for indices $i\in[m]$ and $j\in[k]$. Consider any Boolean-valued function $T:(\mathcal{X}\times\{0,1\})^{m}\times\{0,1\}^{\ell}\to\{0,1\}$ and let the function $\widetilde{T}_{k}:(\mathcal{X}\times\{0,1\})^{m}\to[0,1]$ be
$\widetilde{T}_{k}(x,y)=\biggl[\frac{\gamma}{2}\sum_{j=1}^{k}\sigma_{j}\bm{1}\bigl[\forall i\in[m],\,y_{i}=\bm{1}[f_{j}(x_{i})\geq t_{ij}]\bigr]\biggr]_{0}^{1}.$
Let $\widetilde{T}_{0}=0$. If $f_{j}\in\mathcal{S}_{<(2/\delta^{2}),(\delta/2)}(\mathcal{R}(T)\cup\mathcal{R}(\widetilde{T}_{j-1}))$ for each index $j\in\mathbb{N}$, then there exists a circuit, which we call the classifier circuit for $\widetilde{T}_{k}$, which has the following properties:
  • The classifier receives as input $p=\mathrm{poly}(mk\log(1/\gamma)/\delta)$ Boolean values $r_{j}(x)$ for some restriction functions $r_{1},\ldots,r_{p}\in\mathcal{R}(T)$ evaluated at some particular point $x\in\mathcal{X}$;
  • The classifier uses $q=\mathrm{poly}(mk\log(1/\gamma)/\delta)$ Boolean circuit gates;
  • The classifier outputs the $mk$ Boolean values $\bm{1}[f_{j}(x)\geq t_{ij}]$ for all $i\in[m]$ and $j\in[k]$.
Proof.
We will prove the lemma by induction on $k$. Suppose we have a classifier circuit for $\widetilde{T}_{k-1}$ using $p$ functions $r_{1},\ldots,r_{p}$ from $\mathcal{R}(T)$ and $q$ circuit gates. We will show how to construct a classifier circuit for $\widetilde{T}_{k}$ while increasing $p$ and $q$ by at most $\mathrm{poly}(mk\log(1/\gamma)/\delta)$ each.
By assumption, our circuit already computes the values $\bm{1}[f_{j}(x)\geq t_{ij}]$ for all indices $i\in[m]$ and $j\in[k-1]$. Therefore, we just need to modify the circuit to also compute the values $\bm{1}[f_{k}(x)\geq t_{ik}]$ for each index $i\in[m]$. To this end, recall that
$f_{k}\in\mathcal{S}_{<(2/\delta^{2}),(\delta/2)}\bigl(\mathcal{R}(T)\cup\mathcal{R}(\widetilde{T}_{k-1})\bigr).$
This means that the function $f_{k}$ is a structured sum of at most $2/\delta^{2}$ restrictions of either $T$ or $\widetilde{T}_{k-1}$. We will consider each of these restrictions separately, depending on whether they came from $T$ or $\widetilde{T}_{k-1}$. For the restrictions that came from $T$, add each one to the existing list $r_{1},\ldots,r_{p}$. This increases $p$, the length the list, by at most $2/\delta^{2}$. Next, consider a restriction $r$ obtained by hard-wiring all inputs to $\widetilde{T}_{k-1}$ except for $x_{i}$, for some index $i\in[m]$. Then, $r$ has the form
$r(x)=\biggl[\frac{\gamma}{2}\sum_{j=1}^{k-1}\sigma_{j}\bm{1}\big[y_{i}=\bm{1}[f_{j}(x)\geq t_{ij}]\text{ and }y_{i^{\prime}}=\bm{1}[f_{j}(x_{i^{\prime}})\geq t_{i^{\prime}j}]\text{ for all }i^{\prime}\neq i\bigr]\biggr]_{0}^{1},$
for some fixed sequence $x_{\neq i}$ comprising values $x_{i^{\prime}}$ for all $i^{\prime}\neq i$ and some fixed labels $y\in\{0,1\}^{m}$. Notice that the truth values of the conditions $y_{i^{\prime}}=\bm{1}[f_{j}(x_{i^{\prime}})\geq t_{i^{\prime}j}]$ that appear in the formula for $r$ are fixed functions of $x_{\neq i}$ and $y$. Therefore, they too may be hard-wired in advance. Consequently, the formula for $r$ can be substantially simplified to only depend on some fixed subset of indices $I\subseteq[k-1]$ in the summation:
$r(x)=\biggl[\frac{\gamma}{2}\sum_{j\in I}\sigma_{j}\bm{1}\bigl[y_{i}=\bm{1}[f_{j}(x)\geq t_{ij}]\bigr]\biggr]_{0}^{1}.$
By the inductive hypothesis, our existing classifier circuit has already computed the Boolean values $\bm{1}[f_{j}(x)\geq t_{ij}]$ for all $j\in I$. Therefore, using the formula above, we can compute $r(x)$ using just $\mathrm{poly}(k\log(1/\gamma))$ additional circuit gates. Since $f_{k}$ is a sum of $2/\delta^{2}$ of these restrictions, along with some others from the list $r_{1},\ldots,r_{p}$, we deduce that $f_{k}$ can be computed using just $\mathrm{poly}(mk\log(1/\gamma)/\delta)$ additional circuit gates. Of course, once we have $f_{k}(x)$, the Boolean values $\bm{1}[f_{k}(x)\geq t_{ik}]$ for each index $i\in[m]$ are similarly inexpensive to compute. ∎
We conclude this section with a short proof of Theorem 1.3, the converse to Theorem 1.2.
See 1.3
Proof.
We are given a partition of $\mathcal{X}$ into sets $S_{1},\ldots,S_{k}$, each of which has complexity at most $s$, such that whether or not $f$ belongs to $\mathcal{Q}$ can be determined from its $k$ densities
$\mu_{j}(f)=\operatorname*{\mathbb{E}}\bigl[f(x)\bm{1}[x\in S_{j}]\bigr].$
As usual, the above expectation is over $x\sim\mathcal{D}$. Let $\mu(f)\in[0,1]^{k}$ denote the vector of $k$ densities of $f$, and let $\hat{\mu}_{m}(f)$ denote its empirical estimate given $m$ independent samples $(x_{i},f(x_{i}))$. By Hoeffding’s inequality and a union bound over the $k$ sets, with probability at least $2/3$, we have $\lVert\hat{\mu}_{m}(f)-\mu(f)\rVert_{1}\leq k\delta$ as long as $m=O(\log(k)/\delta^{2})$.
Consider the function $T$ which takes as input a $\delta$-granular density profile $v\in\{0,\delta,\ldots,1-\delta,1\}^{k}$, and outputs $1$ if some $f\in\mathcal{Q}$ has a nearly matching profile, meaning that $\lVert\mu(f)-v\rVert_{1}\leq 2k\delta$. For brevity, let $\pi_{\delta}(v)$ denote the $\delta$-granular coordinate-wise rounding of $v\in[0,1]^{k}$ to multiples of $\delta$.
Fix any $f:\mathcal{X}\to\{0,1\}$. With probability at least $2/3$, we have
$\lVert\mu(f)-\pi_{\delta}(\hat{\mu}_{m}(f))\rVert_{1}\leq\lVert\mu(f)-\hat{\mu}_{m}(f)\rVert_{1}+\lVert\hat{\mu}_{m}(f)-\pi_{\delta}(\hat{\mu}_{m}(f))\rVert_{1}\leq 2k\delta.$
Therefore, $f\in\mathcal{Q}$ implies $T(\pi_{\delta}(\hat{\mu}_{m}(f))=1$. Conversely, if $T(\pi_{\delta}(\hat{\mu}_{m}(f)))=1$, then for some $\tilde{f}\in\mathcal{Q}$,
$\lVert\tilde{f}-f\rVert_{1}=\lVert\mu(\tilde{f})-\mu(f)\rVert_{1}\leq\lVert\mu(\tilde{f})-\pi_{\delta}(\hat{\mu}_{m}(f))\rVert_{1}+\lVert\pi_{\delta}(\hat{\mu}_{m}(f))-\mu(f)\rVert_{1}\leq 4k\delta,$
so $f\in\mathcal{Q}_{4k\delta}$. Setting $\delta=\varepsilon/4k$, we see that running $T$ on the rounded empirical estimate $\pi_{\delta}(\hat{\mu}_{m}(f))$ yields a valid tester for $\mathcal{Q}$ with proximity $\varepsilon$. Since membership in each set $S_{i}$ can be computed with a circuit of size at most $s$, we can determine which parts contain each of the $m=(k/\varepsilon)^{O(1)}$ samples with a circuit of size $(k/\varepsilon)^{O(1)}s$. Maintaining empirical averages can similarly be done with a circuit of size $(k/\varepsilon)^{O(1)}$. Finally, any post-processing function $T:\{0,\delta,\ldots,1-\delta,1\}^{k}\to\{0,1\}$ with $\delta=\Omega(\varepsilon/k)$ can be computed with at most $(k/\varepsilon)^{O(k)}$ additional circuit gates by brute force. ∎
4 Extensions
When combined, Theorems 1.2 and 1.3 give a characterization of properties testable from a constant number of samples. The characterization also holds for the class of properties testable from $m(n)$ samples, where $m(n)=O(\log^{(k)}(n))$ for all $k\in\mathbb{N}$. (Here, $\log^{(k)}$ denotes the $k$-fold iterated logarithm.) In fact, even for properties testable from as many as $O(\log n)$ samples, the more interesting direction of our equivalence (efficient testability implies structured symmetry, or Theorem 1.2) continues to hold.
In this section, we push further in this direction. In Sections 4.1 and 4.2, we study properties that are testable from larger sample sizes than $O(\log n)$. Finally, in Section 4.3, we sketch an argument showing that the results of this paper, although stated in the language of property testing of Boolean functions, can be easily rephrased in the language of testing high-entropy distributions over an arbitrary domain.
4.1 From $O(\log n)$- to $O(n)$-Sample Testability via Consistency Counting
The first extension we consider concerns the number of parts in the partition. It was previously known that $m$-sample testing implies symmetry with $2^{2^{O(m)}}$ parts [BY19]. (Similarly, in Section 3, we showed that efficient $m$-sample testing implies structured symmetry with $2^{2^{O(m)}}$ parts.) In this section, we exponentially improve the dependence on the tester’s sample size, reducing the number of structured sets to just $2^{O(m)}$. Phrased differently, while [BY19] and Theorem 1.2 only give nontrivial conclusions for properties testable with $O(\log n)$ samples, where $2^{n}$ is the domain size, the result of this section remains meaningful for properties testable with $O(n)$ samples.
In order to achieve this improvement, we allow overlaps among the sets that we had previously insisted form a partition. Of course, allowing overlaps requires a corresponding change to the very meaning of “symmetry.” To motivate our new definition, note that a Boolean function property is symmetric with respect to a partition if and only if it depends only on the function’s densities within the parts of the partition. Analogously, in this section, we will show that the property depends only on the function’s densities on a collection of (possibly overlapping) sets. In fact, we will prove an even stronger result: Any $m$-sample testable property essentially boils down to performing consistency counting on a collection of $2^{O(m)}$ special functions.
Definition 4.1 (Consistency Counting).
We say that a collection of $m$ samples $(x_{i},y_{i})$ is consistent with a function $f:\mathcal{X}\to\{0,1\}$ if $y_{i}=f(x_{i})$ for all indices $i\in[m]$. Given a family of good functions $\mathcal{F}_{+}$ and a family of bad functions $\mathcal{F}_{-}$, the $(m,\mathcal{F}_{+},\mathcal{F}_{-})$-consistency counter $T:(\mathcal{X}\times\{0,1\})^{m}\to\{0,1\}$ accepts inputs that are consistent with more good functions than bad.
Clearly, if we run a consistency counter on samples labeled by an unknown function, then the expected output of the tester depends only on the function’s densities on the sets $\{x\in\mathcal{X}\mid f^{\prime}(x)=1\}$, where $f^{\prime}$ is either one of the special “good” functions $\mathcal{F}_{+}$ or one of the special “bad” functions $\mathcal{F}_{-}$. Thus, consistency counting is inherently density-based. However, the converse is not necessarily true. Consistency counters are especially simple density-based testers, which just check whether each of the $k$ special functions could have plausibly labeled the $m$ observed samples. The following theorem, which is the main result of this section, states that every property that is testable from few samples has a consistency counter using a similar number of samples.
Theorem 4.2.
If a property $\mathcal{P}$ is testable with proximity $\varepsilon$ using $m$ samples, then it is also testable with proximity $\varepsilon$ using some $(O(m),\mathcal{F}_{+},\mathcal{F}_{-})$-consistency counter with $\lvert\mathcal{F}_{+}\cup\mathcal{F}_{-}\rvert\leq 2^{O(m)}$.
Note that there is a straightforward converse to Theorem 4.2: By definition, the $(m,\mathcal{F}_{+},\mathcal{F}_{-})$-consistency counter uses just $m$ samples, so if a property is testable with such a consistency counter, then it is testable with $m$ samples.
The proof of Theorem 4.2 is broadly similar to that of Theorem 1.2, but it makes additional use of the additive structure of the simulator from the complexity-theoretic regularity lemma. Since we do not consider computational complexity in this section, we will not require the supersimulators lemma like we did in the previous section. Instead, our proof will only require the basic complexity-theoretic regularity lemma.
Proof of Theorem 4.2.
As in the proof of Theorem 1.2, we are given a property $\mathcal{P}\subseteq\{\mathcal{X}\to\{0,1\}\}$ and a tester $T_{0}:(\mathcal{X}\times\{0,1\})^{m_{0}}\times\{0,1\}^{\ell_{0}}\to\{0,1\}$ that uses $m_{0}$ samples. Since $T_{0}$ is a valid tester for $\mathcal{P}$, it has success probability at least $2/3$. By running $O(1)$ independent copies of $T_{0}$ and taking a majority vote, we can boost the success probability from $2/3$ to $11/12$ (or, for that matter, any constant that is strictly less than $1$). Call the resulting tester $T:(\mathcal{X}\times\{0,1\})^{m}\times\{0,1\}^{\ell}\to\{0,1\}$, where $m=O(m_{0})$ and $\ell=O(\ell_{0})$.
Next, define $x$, $y$, $r$, and $\bar{T}$ as in the proof of Theorem 1.2. In order to construct a consistency counter that is a valid tester for $\mathcal{P}$, consider the distinguisher family $\mathcal{C}\subseteq\{(\mathcal{X}\times\{0,1\})^{m}\to\{0,1\}\}$ comprising all possible $m$-fold consistency functions:
$\mathcal{C}=\Bigl\{(x,y)\mapsto\bm{1}\bigl[y_{1}=f(x_{1})\land\cdots\land y_{m}=f(x_{m})\bigr]\;\Big|\;f:\mathcal{X}\to\{0,1\}\Bigr\}.$
Next, fix $\gamma>0$. By Lemma 2.4 (complexity-theoretic regularity), there exists a function $\widetilde{T}\in\mathcal{S}_{k,\gamma/2}(\mathcal{C})$ that is a $(\mathcal{C},\gamma)$-regular simulator for $\bar{T}$, where $k<2/\gamma^{2}$. Then, by Lemma 3.4 (tester simulation), we have
$\operatorname*{\mathbb{E}}\bigl[T(x,y,r)\bigr]=\operatorname*{\mathbb{E}}\bigl[\bar{T}(x,y)\bigr]\approx_{2^{m}\gamma}\operatorname*{\mathbb{E}}\bigl[\widetilde{T}(x,y)\bigr].$
In other words, replacing $T$ with $\widetilde{T}$ changes the expected output by at most $2^{m}\gamma$, which is strictly less than $1/12$ for appropriately chosen $\gamma=\Theta(1/2^{m})$. Next, by the complexity-theoretic regularity lemma, we know that $\widetilde{T}$ has the form
$\widetilde{T}=\biggl[\frac{\gamma}{2}\bigl(\sigma_{1}F_{1}+\cdots+\sigma_{k}F_{k}\bigr)\biggr]_{0}^{1}$
where each term $F_{j}$ belongs to the family $\mathcal{C}$ and $\sigma_{1},\ldots,\sigma_{k}\in\{\pm 1\}$ are arbitrary signs. By the definition of $\mathcal{C}$, each function $F_{j}$ is testing for consistency with some function $f_{j}:\mathcal{X}\to\{0,1\}$.
Define the set of “good” functions $\mathcal{F}_{+}$ to be the functions $f_{j}$ for which $\sigma_{j}=+1$, and define the set of “bad” functions $\mathcal{F}_{-}$ to be the set of functions $f_{j}$ for which $\sigma_{j}=-1$. Then, the output of $\widetilde{T}$ depends only on the difference between the number of good and bad functions that fit the observed sample:
$\widetilde{T}(x,y)=\biggl[\frac{\gamma}{2}\Bigl\lvert\big\{f\in\mathcal{F}_{+}\;\big|\;\forall i\in[m],\,y_{i}=f(x_{i})\bigr\}\Bigr\rvert-\frac{\gamma}{2}\Bigl\lvert\big\{f\in\mathcal{F}_{-}\;\big|\;\forall i\in[m],\,y_{i}=f(x_{i})\bigr\}\Bigr\rvert\biggr]_{0}^{1}.$
In particular, $\widetilde{T}(x,y)>1/2$ if and only if the $(m,\mathcal{F}_{+},\mathcal{F}_{-})$-consistency counter outputs Accept.
Finally, we show that this consistency counter is a valid tester for $\mathcal{P}$. For this, suppose that $f\in\mathcal{P}$. By assumption, $T$ has success probability at least $11/12$, and the expected output of $\widetilde{T}$ differs from that of $T$ by at most $1/12$. Therefore, the probability that the consistency counter mistakenly outputs Reject is
$\Pr\Bigl[\widetilde{T}(x,y)\leq\frac{1}{2}\Bigr]=\Pr\Bigl[1-\widetilde{T}(x,y)\geq\frac{1}{2}\Bigr]\leq 2\cdot\operatorname*{\mathbb{E}}\bigl[1-\widetilde{T}(x,y)\bigr]\leq 2\cdot\Bigl(\operatorname*{\mathbb{E}}\bigl[1-T(x,y,r)]+\frac{1}{12}\Bigr)\leq\frac{1}{3}.$
Similarly, if $f\notin\mathcal{P}_{\varepsilon}$, then the consistency counter mistakenly outputs Accept with probability at most $1/3$. We conclude that the $(m,\mathcal{F}_{+},\mathcal{F}_{-})$-consistency counter is a valid tester for $\mathcal{P}$. ∎
4.2 Beyond $O(n)$-Sample Testability via Regularity Templates
The second extension we consider concerns regularity templates—a concept we define shortly. In Theorem 1.2, we studied properties that are efficiently testable from $O(\log n)$ samples, and in Theorem 4.2, we studied properties that are testable from $O(n)$ samples. What about properties that require larger sample sizes to test? In this section, we show that if a property is testable using $m$ samples and a circuit of size $s$, then the property is essentially testing for compatibility with a collection of regularity templates, each of complexity $O(m^{2}s)$. In particular, this result is nontrivial for all arbitrary (subexponential) sample sizes $m$ and circuit sizes $s$.
We are motivated by the classic characterization of testable graph properties, due to Alon et al. [AFNS06]. In that work, it was shown that a graph property is testable if and only if the property is regular-reducible, meaning that it essentially amounts to checking compatibility with a bounded collection of regularity templates. In the context of graph regularity, a regularity template is represented as a list of $\binom{k}{2}$ pairwise densities representing the densities between a partition of the graph into $k$ blocks. In the context of efficiently testing Boolean functions, we shall model a “regularity template” by a regular simulator for the function to be tested. We shall define the property of being “comptaible” with a regularity template via the usual notion of indistinguishability, as per Definition 2.2:
Definition 4.3.
Given a collection $\mathcal{T}\subseteq\{\{0,1\}^{n}\to[0,1]\}$ of functions, define the compatibility property $\mathcal{T}_{s,\delta}$ to be the set of functions that are $(s,\delta)$-indistinguishable from some function in $\mathcal{T}$.
The following result states that small-circuit testability (even for a moderate or large number of samples) reduces to testing compatibility with a size-bounded collection of regularity templates.
Theorem 4.4.
If the property $\mathcal{P}$ is testable with proximity $\varepsilon$ using $m$ samples and a circuit of size $s$, then there exists a set $\mathcal{T}$ of template circuits, each of size $O(m^{2}s)$, such that $\mathcal{P}\subseteq\mathcal{T}_{s,\frac{1}{13m}}\subseteq\mathcal{P}_{\varepsilon}$.
Proof.
We are given a property $\mathcal{P}$ and a valid $m$-sample, size-$s$ tester $T$. Set $\delta=1/13m$. By Lemma 2.4 (complexity-theoretic regularity), every Boolean function has an $(s,\delta)$-regular simulator computable by a circuit of size $O(m^{2}s)$. With this in mind, let $\mathcal{T}$ be the set of $(s,\delta)$-regular simulators of size $O(m^{2}s)$ for functions in $\mathcal{P}$. Clearly, $\mathcal{P}\subseteq\mathcal{T}_{s,\delta}$.
Conversely, we claim that $\mathcal{T}_{s,\delta}\subseteq\mathcal{P}_{\varepsilon}$. To prove this, suppose for the sake of contradiction that there exists a function $f\in\mathcal{T}_{s,\delta}\setminus\mathcal{P}_{\varepsilon}$. Let $\tilde{f}$ be a function in $\mathcal{T}$ that is $(s,\delta)$-indistinguishable from $f$, and let $f^{\prime}$ be a function in $\mathcal{P}$ that is $(s,\delta)$-indistinguishable from $\tilde{f}$. Since $T$ has a circuit of size $s$, so does every function in the class $\mathcal{R}(T)$ of restrictions of $T$. Thus, $(s,\delta)$-indistinguishability implies $(\mathcal{R}(T),\delta)$-indistinguishability. Therefore, letting $\bar{T}$ denote the expectation of $T$ over its internal randomness as usual, Lemma 3.2 (oracle simulation) implies
$\operatorname*{\mathbb{E}}\bigl[T(x,f(x),r)\bigr]\approx_{2m\delta}\operatorname*{\mathbb{E}}\bigl[\bar{T}(x,\tilde{y})\bigr]\approx_{2m\delta}\operatorname*{\mathbb{E}}\bigl[T(x,f^{\prime}(x),r)\bigr]\geq\frac{2}{3},$
where the expectation is over independent $x_{i}\sim\{0,1\}^{n}$ and $\tilde{y}_{i}|x_{i}\sim\mathcal{B}(\tilde{f}(x))$ and random seed $r$. In other words, replacing $f$ with $f^{\prime}$ changes the probability that $T$ outputs Accept by at most $4m\delta$, which is strictly less than $1/3$ for our choice of $\delta$. However, since $f\notin\mathcal{P}_{\varepsilon}$, we also have $\operatorname*{\mathbb{E}}[T(x,f(x),r)]\leq 1/3$, which is a contradiction. We conclude that $\mathcal{P}\subseteq\mathcal{T}_{s,\delta}\subseteq\mathcal{P}_{\varepsilon}$. ∎
We conclude this section with a partial converse to Theorem 4.4. Note that it is not a full converse because it sacrifices computational efficiency. In other words, the tester that is provides is not necessarily computable with a small circuit. The main idea behind the proof of the converse is a trick involving convex combinations, which was used in [AFNS06] to prove that compatibility with a fixed regularity template is testable. The rest of the proof is a series of standard applications of Chernoff/Hoeffding bounds.
Theorem 4.5.
Let $s\leq t$ with $t\log(t)\leq O(\varepsilon^{2}\delta^{2}2^{n})$. If $\mathcal{T}$ is a set of template circuits, each of size at most $t$, then the property $\mathcal{T}_{s,\delta}$ is testable with proximity $\varepsilon$ from $O(t\log(t)/\varepsilon^{2}\delta^{2})$ samples.
Proof.
Let $g$ denote the function we wish to test. For each $h\in\mathcal{T}$, size-$s$ $f$, and $\sigma\in\{\pm 1\}$, consider
$\operatorname*{\mathbb{E}}\Bigl[\sigma f(x)\bigl(g(x)-h(x)\bigr)\Bigr].$
By definition, $g$ has the property $\mathcal{T}_{s,\delta}$ if and only if there exists some template $h\in\mathcal{T}$ such that the above quantity is at most $\delta$ for all size-$s$ $f$ and $\sigma\in\{\pm 1\}$. For now, consider a fixed $h\in\mathcal{T}$. Note that there are at most $2^{O(s\log s)}$ circuits $f$ of size at most $s$. Therefore, by Hoeffding’s inequality, we can estimate the quantity displayed above for all size-$s$ functions $f$ and signs $\sigma$ up to error $\alpha$ with failure probability $\beta$, given $m=O((s\log(s)+\log(1/\beta))/\alpha^{2})$ labeled samples. Let us output Accept if all of these estimates are at most $\delta+\alpha$, and output Reject otherwise. Clearly, if $g\in\mathcal{T}_{s,\delta}$, then the tester accepts with probability with at least $1-\beta$.
Conversely, we claim that with probability at least $1-\beta$, if the tester accepts, then $g$ is $\varepsilon$-close to $\mathcal{T}_{s,\delta}$. To prove this, first note that $g$ is $(s,\delta+\alpha)$-indistinguishable from $h$. Consequently, any convex combination $\lambda h+(1-\lambda)g$ is $(s,(1-\lambda)(\delta+\alpha))$-indistinguishable from $h$. By performing randomized rounding on the output of this combination, we obtain a Boolean-valued function $\tilde{g}$ that is $2\lambda$-close to $g$ with probability at least $1-\beta$, for a sufficiently large domain size $\lvert\mathcal{X}\rvert=O(\log(1/\beta))$. Moreover, this $\tilde{g}$ is $(s,(1-\lambda)(\delta+\alpha)+\gamma)$-indistinguishable from $h$ with probability at least $1-\beta$, for $\lvert\mathcal{X}\rvert=O((s\log(s)+\log(1/\beta))/\gamma^{2})$. If we set $\lambda=(\alpha+\gamma)/(\delta+\alpha)$ and $\gamma\leq\alpha\leq\varepsilon\delta/4$, then $\lambda\leq\varepsilon/2$, in which case $\tilde{g}$ is both $\varepsilon$-close to $g$ and $(s,\delta)$-indistinguishable from $h$, as desired.
Until now, we have considered a fixed template $h\in\mathcal{T}$ and used only $O((s\log(s)+\log(1/\beta))/\varepsilon^{2}\delta^{2})$ samples. However, since every circuit in $\mathcal{T}$ has size at most $t$, we can run our tester for all $h\in\mathcal{T}$, at the cost of a $2^{O(t\log t)}$ factor blowup in the failure probability $\beta$. Since $s\leq t$, we conclude that $\mathcal{T}_{s,\delta}$ is testable with proximity $\varepsilon$ from $O(t\log(t)/\varepsilon^{2}\delta^{2})$ samples. ∎
4.3 From Function Testing to High-Entropy Distribution Testing
In this section, we sketch the third and final extension we consider, which concerns high-entropy, or dense, distribution testing. So far, we have focused on the problem of testing properties of Boolean functions from samples, but in some sense, this framing is too restrictive. Indeed, in this section, we show that our main result can be easily extended to testing properties of distributions on arbitrary domains, provided that the distribution to be tested is known to have high density (equivalently, high min-entropy).
First, we review the relevant notions of density and entropy. Given a finite domain $\mathcal{X}$, fix a distribution $\mathcal{D}_{0}\in\Delta(\mathcal{X})$, which we call the base distribution or reference distribution. For example, when $\mathcal{X}=\{0,1\}^{n}$, we take $\mathcal{D}_{0}$ to be the uniform distribution on $\mathcal{X}$. Given another distribution $\mathcal{D}$ whose support is contained in that of $\mathcal{D}_{0}$, we say that $\mathcal{D}$ is $\mu$-dense in $\mathcal{D}_{0}$ if the ratio of $\mathcal{D}$ to $\mathcal{D}_{0}$ is at most $1/\mu$ pointwise. For example, when $\mathcal{X}=\{0,1\}^{n}$, a distribution is $\mu$-dense if it is a uniform distribution over a set of size $\mu\lvert\mathcal{X}\rvert$, or a convex combination thereof. In this case, we equivalently say that the distribution has min-entropy at least $k=n-\log(1/\mu)$.
For another example, suppose that we sample $x\sim\{0,1\}^{n}$ uniformly, and then set $y=f(x)$ for some fixed Boolean function $f$. Then, the joint distribution of the pair $(x,y)$ is $\mu$-dense in the uniform distribution over $\{0,1\}^{n+1}$ for $\mu=1/2$. Indeed, this is precisely the fact that we used in the proof of Lemma 3.4, which lead to the $(1/\mu)^{m}=2^{m}$ blowup in the regularity parameter $\gamma$. In fact, a lower bound on the density of the distribution of $(x,y)$ was essentially the only thing our proof needed; it was not important for $x$ to be uniform and $y$ to be a deterministic function of $x$.
Dense Distribution Testing
With the aforementioned example in mind, we now consider a natural analogue of the property testing framework from Section 2 that applies not just to uniform $n$-bit strings with Boolean labels, but also more general dense distribution properties. For this, recall that the total variation distance between two distributions $\mathcal{D}_{0},\mathcal{D}_{1}\in\Delta(\mathcal{X})$ is their maximum absolute difference over any subset of the domain: $\max_{S\subseteq\mathcal{X}}\,\lvert\mathcal{D}_{0}(S)-\mathcal{D}_{1}(S)\rvert$.
Definition 4.6 (Densely Testable Property).
Let $\mathcal{P}\subseteq\Delta(\{0,1\}^{n})$ be a distribution property. We say that $\mathcal{P}$ is $\mu$-densely testable with proximity parameter $\varepsilon>0$ if there exists a randomized circuit $T$ of size $s$ that receives as input $m$ samples $x_{i}\sim\mathcal{D}$ for some $\mathcal{D}\in\Delta(\{0,1\}^{n})$, always outputs either Accept or Reject, and meets the following two requirements:
  • If $\mathcal{D}\in\mathcal{P}$ and $\mathcal{D}$ is $\mu$-dense, then $T$ outputs Accept with probability at least $2/3$.
  • If $\mathcal{D}$ is $\varepsilon$-far from $\mathcal{P}$ and $\mathcal{D}$ is $\mu$-dense, then $T$ outputs Reject with probability at least $2/3$.
In each condition, the probability is computed over randomness in the sample and internal to $T$.
In our to state our generalized main result for dense distribution testing, we will again need to invoke the notion of “structured symmetry” that we have been using throughout this paper. However, we emphasize that in the context of distribution testing, this is not the same as the more standard notion of label-invariance. Rather, we say that a distribution property $\mathcal{P}$ is $k$-part symmetric if there is a partition of $\{0,1\}^{n}$ into disjoint parts $S_{1},\ldots,S_{k}$ such that $\mathcal{P}$ is invariant under any redistribution of probability mass within each part. (More formally, whether or not $\mathcal{D}\in\mathcal{P}$ should depend only on the $k$ densities $\mathcal{D}(S_{i})$ for $i\in[k]$.)
We now state the main result of this section, which directly generalizes Theorem 1.2 (efficient testing of Boolean functions implies structured symmetry) to the setting of $\mu$-dense distribution testing. The converse, of course, is straightforward. Since the proof of Theorem 4.7 is extremely similar to that of Theorem 1.2, which we carried out in detail in Section 3, we simply sketch the main alterations to the argument.
Theorem 4.7.
Fix $\mu\in(0,1/2)$. If a distribution property $\mathcal{P}$ is $\mu$-densely testable with proximity $\varepsilon$ using $m$ samples and a circuit of size $s$, then $\mathcal{P}\subseteq\mathcal{Q}\subseteq\mathcal{P}_{\varepsilon}$ for some $2^{(1/\mu)^{O(m)}}$-part symmetric distribution property $\mathcal{Q}$ with partition complexity $(1/\mu)^{O(m)}s$.
Proof Sketch.
First, some notation. Given $f:\{0,1\}^{n}\to[0,1/\mu]$, let $\mathcal{D}_{f}$ denote the $\mu$-dense distribution over $\{0,1\}^{n}$ with mass function $f(x)2^{-n}$. Given $x\in(\{0,1\}^{n})^{m}$, let $f^{(m)}(x)=\prod_{i=1}^{m}f(x_{i})$.
Our strategy will be to generalize Lemma 3.2 (oracle simulation) and Lemma 3.4 (tester simulation), which were the key building blocks of Theorem 1.2. To generalize the oracle simulation lemma, consider a tester $T:(\{0,1\}^{n})^{m}\times\{0,1\}^{\ell}\to\{0,1\}$ computable by a circuit of size $s$. If $\mu\tilde{f}$ is $(s,\delta)$-indistinguishable from $\mu f$, then by the same hybrid argument as before,
$\mathbb{E}_{f}\bigl[T(x,r)\bigr]=\frac{1}{\mu}\operatorname*{\mathbb{E}}\bigl[\mu f^{(m)}(x)T(x,r)]\approx_{\frac{m\delta}{\mu}}\frac{1}{\mu}\operatorname*{\mathbb{E}}\bigl[\mu\tilde{f}^{(m)}(x)T(x,r)\bigr]=\mathbb{E}_{\tilde{f}}\bigl[T(x,r)\bigr],$
where $\mathbb{E}_{f}$ denotes the expectation over $x\sim\mathcal{D}_{f}^{m}$ and $\operatorname*{\mathbb{E}}$ denotes the expectation over uniform $x$. Next, to generalize the tester simulation lemma, suppose that $\widetilde{T}$ and $\bar{T}$ (the mean function of $T$) are $\gamma$-indistinguishable with respect to $m$-fold thresholds of $\tilde{f}$. Then, for a uniform $t\sim[0,1]^{m}$,
$\mathbb{E}_{\tilde{f}}\bigl[\bar{T}(x)\bigr]=\mu^{-m}\operatorname*{\mathbb{E}}\bigl[\bm{1}[\mu\tilde{f}(x)\geq t]\,\bar{T}(x)\bigr]\approx_{\mu^{-m}\gamma}\mu^{-m}\operatorname*{\mathbb{E}}\bigl[\bm{1}[\mu\tilde{f}(x)\geq t]\,\widetilde{T}(x)\bigr]=\mathbb{E}_{\tilde{f}}\bigl[\widetilde{T}(x)\bigr].$
Having generalized the two lemmas, we proceed as in the proof of Theorem 1.2. Set $\delta=\Theta(\mu/m)$ and $\gamma=\Theta(\mu^{m})$. By Lemma 2.6 (supersimulators), there exists a circuit $\widetilde{T}$ of size $\tilde{s}$ that fools all $m$-fold thresholds of functions $\tilde{f}$ computable by a circuit of size at most $O(\max(s,\tilde{s})/\delta^{2})$. Next, by Lemma 2.4 (complexity-theoretic regularity), given any $\mu$-dense distribution $\mathcal{D}_{f}$, the function $\mu f$ is $(\max(s,\tilde{s}),\delta)$-indistinguishable from some $\mu\tilde{f}$ computable in size $O(\max(s,\tilde{s})/\delta^{2})$. Thus, combining the two generalized lemmas, we have that, say, $\lvert\mathbb{E}_{f}[T(x,r)]-\mathbb{E}_{f}[\widetilde{T}(x)]\rvert<1/6$. This implies that $\mathcal{P}\subseteq\mathcal{Q}\subseteq\mathcal{P}_{\varepsilon}$, where $\mathcal{Q}$ is the property that the expected output of $\widetilde{T}$ is at least $1/2$. As before, the structured symmetry can be read off from the structure of the circuit $\widetilde{T}$ provided by the supersimulators lemma. The bounds $2^{(1/\mu)^{O(m)}}$ and $(1/\mu)^{O(m)}s$ on the number of parts and the partition complexity, respectively, follow from the same counting arguments as before with our new choices of the parameters $\delta=\Theta(\mu/m)$ and $\gamma=\Theta(\mu^{m})$. ∎
Appendix A Simple Simulation with One Projection
In this section, we give a short, unified proof of Lemmas 2.4 and 2.6, which are the specific versions of the complexity-theoretic regularity lemma [TTV09] and supersimulators lemma [DT25] on which our work relies. The version of the supersimulators lemma in [DT25] produced an approximation to the target function $g$ with a nested sequence of projections onto the unit interval, i.e.
$h(x)=\biggl[\cdots\Bigl[\bigl[[\delta f_{1}(x)]_{0}^{1}+\delta f_{2}(x)\bigr]_{0}^{1}+\cdots\Bigr]_{0}^{1}+\delta f_{k}(x)\biggr]_{0}^{1},$
for appropriate distinguisher functions $f_{j}$. The same is true of some proofs of the complexity-theoretic regularity lemma (e.g. the one from [HKRR18], as observed by [CDV24]). In contrast, Lemmas 2.4 and 2.6 use just a single such projection:
$h(x)=\Bigl[\delta f_{1}(x)+\cdots+\delta f_{k}(x)\Big]_{0}^{1}.$
Interestingly, the proof of the original complexity-theoretic regularity lemma from [TTV09] required a careful case analysis to get by with just this single projection. Our proof in this section is a bit simpler, relying on only the following lemma about prefix sums.
Lemma A.1 (Prefix Sums).
If $a_{1},\ldots,a_{k}\in\mathbb{R}$ and $s_{j}=[a_{1}+\cdots+a_{j}]_{0}^{1}$ and $b\in[0,1]$, then
$\sum_{j=1}^{k}a_{j}(b-s_{j})\leq\frac{1}{2}b^{2}.$
Proof.
We induct on $k$. The base case ($k=0$) is trivial. For $k\geq 1$, we observe that
$\sum_{j=1}^{k}a_{j}(b-s_{j})=\sum_{j=1}^{k-1}a_{j}(s_{k}-s_{j})+\sum_{j=1}^{k}a_{j}(b-s_{k}).$
By the inductive hypothesis, the first sum on the right side is at most $s_{k}^{2}/2$. To conclude, we must show that the second sum on the right side is at most $(b^{2}-s_{k}^{2})/2$. Some algebra does the trick:
$\sum_{j=1}^{k}a_{j}(b-s_{k})=\frac{1}{2}(b^{2}-s_{k}^{2})+\frac{1}{2}\biggl[\Bigl(s_{k}-\sum_{j=1}^{k}a_{j}\Bigr)^{2}-\Bigl(b-\sum_{j=1}^{k}a_{j}\Bigr)^{2}\biggr],$
where the term in square brackets is $\leq 0$ by the definition of $s_{k}$ and the assumption $b\in[0,1]$. ∎
Equipped with this result on prefix sums, we are now ready to prove Lemma 2.6.
See 2.6
Proof.
We will inductively define functions $h_{0},\ldots h_{k-1}:\mathcal{X}\to[0,1]$ and then argue that some $h_{j}$ must be a $(\mathcal{G}(h_{j}),\delta)$-regular simulator for $g$. First, set $h_{0}(x)=0$. For $j\geq 1$, suppose that $h_{j-1}$ is not $(\mathcal{G}(h_{j-1}),\delta)$-regular, and let $f_{j}$ be any function in $\pm\mathcal{G}(h_{j-1})$ such that
$\operatorname*{\mathbb{E}}\Bigl[f_{j}(x)\bigl(g(x)-h_{j-1}(x)\bigr)\Bigr]>\delta,$
where the expectation is over $x\sim\mathcal{D}$. Next, fix $\eta>0$ and define
$h_{j}(x)=\Bigl[\eta\sum_{i=1}^{j}f_{i}(x)\Bigr]_{0}^{1}.$
Since $\lvert f_{j}(x)\rvert\leq 1$, the functions $h_{j}$ and $h_{j-1}$ differ pointwise by at most $\eta$. Thus,
$\operatorname*{\mathbb{E}}\Bigl[f_{j}(x)\bigl(g(x)-h_{j}(x)\bigr)\Bigr]>\delta-\eta.$
Summing over $j\in[k]$ and applying Lemma A.1 with $a_{j}=\eta f_{j}(x)$ and $b=g(x)$ yields
$k\cdot\eta(\delta-\eta)<\operatorname*{\mathbb{E}}\Bigl[\sum_{j=1}^{k}\eta f_{j}(x)\bigl(g(x)-h_{j}(x)\bigr)\Bigr]\leq\frac{1}{2}\operatorname*{\mathbb{E}}\Bigl[g(x)^{2}\Bigr]\leq\frac{1}{2}.$
Setting $\eta=\delta/2$ and $k=2/\delta^{2}$ yields a contradiction. We conclude that one of the functions $h_{j}$ for $0\leq j<k$ must have been an $(\mathcal{G}(h_{j}),\delta)$-regular simulator for $g$, as desired. ∎
We remark that our proof is essentially a special case of the analysis of standard algorithms in online learning and convex optimization, such as Follow-the-Regularized-Leader (FTRL) and mirror descent with lazy projections. Rather than invoking the full strength of this machinery, we have chosen to isolate the minimal technical tool required for the job in Lemma A.1. For more on online convex optimization, we refer the reader to the surveys [Bub15, McM17].